OpenStudy (anonymous):

A bag has 1 red marble, 4 blue marbles, and 3 green marbles. Peter draws a marble randomly from the bag, replaces it, and then draws another marble randomly. What is the probability of drawing 2 blue marbles in a row? Explain your answer.

OpenStudy (anonymous):

@Vocaloid

Mehek (mehek14):

there is 1 red marble and a total of 8 marbles \(\dfrac{1}{8}*\dfrac{1}{8}=\dfrac{1}{64}\)

OpenStudy (anonymous):

OK thank you!

OpenStudy (anonymous):

wait is that all? @Mehek14

Mehek (mehek14):

yup

OpenStudy (anonymous):

k thanks!